OpenStudy (anonymous):

Factor: 9x^4-18x^2+8

OpenStudy (anonymous):

heeelllp

OpenStudy (anonymous):

Let u=x^2 \[9u^2-18u+8\] Factorise like a normal quadratic.

OpenStudy (anonymous):

But it's ^4 I don't know how to factor that

OpenStudy (anonymous):

@bigworm Read what I wrote mate. Pay attention.

OpenStudy (anonymous):

Well mate that didn't help

OpenStudy (anonymous):

That does help @bigworm I told you to factorise this: \[9u^2-18u+8\]

OpenStudy (anonymous):

You let u=x^2 which basically means, everytime you see x^2 you substitute u. And for x^4 it is the same as \[x^4=x^2\times x^2\] \[x^4=u\times u\] \[x^4=u^2\] So now when you see x^4, you substitute u^2

OpenStudy (anonymous):

@bigworm you there?

OpenStudy (anonymous):

or are you waiting for someone to hand over the answer to you?

OpenStudy (anonymous):

I was waiting for someone to solve it out for me so I know how to do the rest of the problems

OpenStudy (anonymous):

Yeah that's a pretty good plan if you want to get low marks for maths.

OpenStudy (anonymous):

In exams, markers don't just want answers. They want to see if you undertood your maths through the working you provided when answering questions.

OpenStudy (anonymous):

understood*